# Kac-Ward solution of the 2D classical and 1D quantum Ising models

### v1

@@ Abstract.
We give a rigorous derivation of the free energy of (i) the classical Ising model on the triangular lattice with translation-invariant coupling constants, and (ii) the one-dimensional quantum Ising model. We use the method of Kac and Ward. The novel aspect is that the coupling constants may have negative signs. We describe the logarithmic singularity of the specific heat of the classical model and the validity of the Cimasoni–Duminil-Copin-Li formula for the critical temperature. We also discuss the quantum phase transition of the quantum model.
@@ 1. Introduction
After the work of Onsager and Kaufman, people found two alternate approaches: combinatorial and fermionic. The former was proposed in 1952 by Kac and Ward [12]; it was later extended by Kasteleyn who noted the connection with dimer systems [14]. Potts [22] and Stephenson [24] used the Kac-Ward method on the triangular lattice, for the free energy and for correlation functions. The fermionic method was proposed in 1964 by Schultz, Mattis, and Lieb [23].
@@
The main goal of this article is to extend the Kac-Ward method to the case of (translation-invariant) coupling constants of arbitrary signs. We work on the triangular lattice, which is the simplest case of frustrated systems with translation-invariant coupling constants. We start with the Cimasoni extension of the Kac-Ward formula for “faithful projections” of non-planar graphs [5] (see also Aizenman and Warzel [2] for a clear exposition). We use it for the torus $\{1,\dots,L\}_{\rm per}\times\{1,\dots,M\}_{\rm per}$ with periodic boundary conditions. The main difficulties involve the non-planarity of the graph. We prove that these difficulties vanish in the limit $L\to\infty$ for fixed $M$. Then we can use the Fourier transform and we obtain the free energy formula for the infinite cylinder ×${\mathbb{Z}}\times\{1,\dots,M\}_{\rm per}$. The Onsager-Houtappel formula immediately follows by taking the limit $M\to\infty$.
@@
Our result for cylinders allows us to consider the one-dimensional quantum Ising model, whose free energy was first calculated in 1970 by Pfeuty [21]. It can be mapped to a 2D classical Ising model in the limit where the extra dimension becomes continuous. We also discuss the occurrence of a “quantum phase transition”.
@@ 2.1. The free energy
In the limit $M\to\infty$, the cylinder free energy density converges to the two-dimensional free energy density:
@@ 2.2. Logarithmic singularity of the specific heat
We are looking for values of $\beta$ where ${\mathsf{f}}$ is not analytic. We show the well-known fact that the second derivative of ${\mathsf{f}}$ (which is related to the physical quantity called the specific heat) has logarithmic singularity at a special value $\beta_{\rm c}$, called the critical point. By Theorem 2.1 (b) we have
@@
It is not hard to verify that the second condition holds true when $\sinh(2\beta_{\rm c}J_{i})+2\sinh(2\beta_{\rm c}J_{3})>0$, $i=1,2$.
@@
We already know that ${\mathsf{f}}(\beta)$ is concave, and therefore continuous. For $\beta\neq\beta_{\rm c}$, we have
@@
We have now verified that the only divergent term in (2.18) is the first one, and divergence is logarithmic indeed. ∎
@@ 2.3. Case $J_{1}=J_{2}$
    If $J_{3}>-1$, there is a unique $\beta_{\rm c}$ such that ${\mathsf{f}}(\beta)$ is analytic in +${\mathbb{R}}_{+}\setminus\{\beta_{\rm c}\}$ and ${\mathsf{f}}^{\prime\prime}(\beta)$ has logarithmic divergence at $\beta_{\rm c}$.
@@
The relation with the free energy is that
@@
Then $g(\beta)$ vanishes precisely when $a(\beta)$ does.
We check in the proof below that the Cimasoni–Duminil-Copin–Li formula holds whenever $J_{1}=J_{2}\geq 0$, and $J_{3}\in{\mathbb{R}}$ is allowed to be negative. One can also check that it does not hold if $J_{1}=J_{2}$ change signs; indeed, the free energy is the same due to symmetries, but the formula is different and has different solutions.
@@ 3. The Kac-Ward identity
We rely on the extension of the Kac-Ward identity for “faithful projections” of non-planar graphs. It was proposed by Cimasoni [5] and used in [13, 2]. In order to accommodate negative weights we need two faithful projections for ${\mathbb{T}}_{L,M}$ with edges between nearest-neighbours. The graphs are $G_{1}$ and $G_{2}$ and they are illustrated in Fig. 2. Here is a full description of the left graph:
@@
  • There are edges represented by “handles” (continuous curves with winding number $-1$) between $(L,j)$ and $(1,j)$ for $1\leq j\leq M$; between $(L,j)$ and $(1,j+1)$ for $11\leq M-$ 1; between $(i,M)$ and $(i,1)$ for $1\leq i\leq L$; between $(i,M)$ and $(i+1,1)$ for $11\leq i\leq L-$ 1.
@@
It is worth noting that the right side of (3.3) is a multi-monomial in $(W_{e,e})$, something that is not apparent in the left side — there are remarkable cancellations indeed. This allows [2] to prove the identity for small $(W_{e,e})$; the extension to larger values is automatic. The determinant cannot be negative and the sign of the square root cannot change.
We define the matrix $K^{\scriptscriptstyle(2)}$ as in (3.1) but $\measuredangle_{2}(e,e^{\prime})$ and $\measuredangle_{2}(e)$ are the corresponding angles on the the faithfull projection $G_{2}$. Analogously, we have $n^{\scriptscriptstyle(2)}_{0}(\Gamma)$ for this projection.
The connection with the Ising model is through the high-temperature expansion, see e.g. [7, Section 3.7.3]. The partition function (2.4) is equal to
@@
$\sqrt{\det(1-K^{\scriptscriptstyle(1)}W)}+\sqrt{\det(1-K^{\scriptscriptstyle(2)}W)}=2\widetilde{Z}_{L,M}\Bigl{(}1-\big{\langle}1_{n_{\rm h}(\Gamma)\,\rm{odd}}\big{\rangle}_{L,M}\Bigr{)}$
@@
We condition on the horizontal handles (including possibly the self-crossing ones). We denote by ${\mathfrak{h}}$ the set of horizontal handles
@@
Then we define the support ${\rm{supp\,}}_{1}{\mathfrak{h}}$, resp.  ${\rm{supp\,}}_{L}{\mathfrak{h}}$, to be the set of vertices of the form $(1,j_{i})$, resp.  $(L,j_{i}^{\prime})$, that appear an odd number of times in ${\mathfrak{h}}$. Let ${\rm{supp\,}}{\mathfrak{h}}={\rm{supp\,}}_{1}{\mathfrak{h}}\cup{\rm{supp\,}}_{L}{\mathfrak{h}}$. We let ~$\tilde{\mathcal{E}}_{L,M}$ be the set of edges of the cylinder (not the torus) $\{1,\dots,L\}\times{\mathbb{T}}_{M}$. We have
$\begin{split}\big{\langle}1_{n_{\rm h}(\Gamma)\,\rm{odd}}\big{\rangle}_{L,M}&=\sum_{|{\mathfrak{h}}|\;{\rm odd}}\langle 1_{{\mathfrak{h}}\subset\Gamma}\rangle_{L,M}\\ &=\sum_{|{\mathfrak{h}}|\;{\rm odd}}\biggl{(}\prod_{i=1}^{k}\tanh J_{(1,j_{i}),(L,j_{i}^{\prime})}\biggr{)}\frac{1}{\tilde{Z}_{L,M}}\sum_{\Gamma\subset\tilde{\mathcal{E}}_{L,M}:\partial\Gamma={\rm{supp\,}}{\mathfrak{h}}}w(\Gamma).\end{split}$ (3.9)
@@
One can expand the determinants as products of directed loops as in [2, Theorem 3.2]. If $\gamma=(e_{1},\dots,e_{k})$ is a directed loop, one can verify that
$\prod_{i=1}^{k}K^{\scriptscriptstyle(1)}_{e,e^{\prime}}=(-1)^{\text{winding}^{\scriptscriptstyle(1)}(\gamma)},\qquad\prod_{i=1}^{k}\widetilde{W}^{\scriptscriptstyle(1)}_{e,e}=(-1)^{\text{winding}^{\scriptscriptstyle(1)}(\gamma)}\prod_{i=1}^{k}W_{e,e}.$ (3.17)
where $\text{winding}^{\scriptscriptstyle(1)}(\gamma)$ is the winding number of the loop $\gamma$ in the projection $G_{1}$. Then each loop gives the same contribution in $\det(1-K^{\scriptscriptstyle(1)}W)$ and $\det(1-\widetilde{K}\widetilde{W}^{\scriptscriptstyle(1)})$. The argument for $\det(1-K^{\scriptscriptstyle(2)}W)$ is the same, with the projection $G_{2}$. ∎
@@
where $t_{1}=\tanh{J_{1}}$, $t_{2}=\tanh{J_{2}}$ and $t_{3}=\tanh{J_{3}}$. Then
@@
(a) By Eq. (3.3) and Lemma 3.3, we obtain that both square roots of the above determinants are real. Further, these determinants are real by Lemma 3.5. (b) This is a consequence of Lemma 3.4; taking the logarithm we obtain Riemann sums. ∎
@@ 4. The 1D quantum Ising model
One application of the cylinder formula of Theorem 2.1 (a) is for the one-dimensional quantum Ising model. It is well-known that it can be mapped to a classical model in $1+1$ dimensions, the extra dimension being continuous. Its free energy was first computed by Pfeuty [21] using the fermionic method of [23]. The results of this section are not new, but the Kac-Ward approach may have more appeal for some readers.
@@
$\begin{split}{\operatorname{Tr\,}}\,{\rm e}^{-\beta H_{L}}\,&=\lim_{n\to\infty}{\operatorname{Tr\,}}\biggl{(}\,{\rm e}^{\frac{\beta}{n}\sum_{i=1}^{L}S_{i}^{\scriptscriptstyle(3)}S_{i+1}^{\scriptscriptstyle(3)}}\,\prod_{i=1}^{L}\bigl{(}1+\tfrac{\beta h}{n}S_{i}^{\scriptscriptstyle(1)}\bigr{)}\biggr{)}^{n}\\ &=\lim_{n\to\infty}\sum_{{\boldsymbol{\sigma}}^{\scriptscriptstyle(1)},\dots,{\boldsymbol{\sigma}}^{\scriptscriptstyle(n)}}\exp\biggl{\{}\frac{\beta}{4n}\sum_{i=1}^{L}\sum_{k=1}^{n}\sigma_{i}^{\scriptscriptstyle(k)}\sigma_{i+1}^{\scriptscriptstyle(k)}\biggr{\}}\prod_{i=1}^{L}\prod_{k=1}^{n}\langle\sigma_{i}^{\scriptscriptstyle(k)}|\bigl{(}1+\tfrac{\beta h}{n}S^{\scriptscriptstyle(1)}\bigr{)}|\sigma_{i}^{\scriptscriptstyle(k+1)}\rangle.\end{split}$ (4.5)
@@
    For fixed $n$ the limit $L\to\infty$ of $f_{L,n}^{\rm qu}(\beta,h)$ exists (and is denoted $f_{\infty,n}^{\rm qu}(\beta,h)$.
@@
The first equality in (b) is clear. For the second equality we use the following estimates, which again follows from estimates on the classical partition function:
@@
$f_{L,n}^{\rm qu}(\beta,h)+\tfrac{\beta}{2L}\geq f_{\infty,n}^{\rm qu}(\beta,h)\geq f_{L,n}^{\rm qu}(\beta,h)-\tfrac{\beta}{2L}.$ (4.9)
@@
This holds for any ${\varepsilon}>0$ provided $n$ is large enough. This proves the second expression in (b). ∎
@@
It can be obtained by taking the limit $J_{1}\to 0$ in Theorem 2.1 (a), as the expression converges to the free energy of the 1D Ising model in ${\mathbb{T}}_{M}$. The latter is easily calculated with the 1D transfer matrices, yielding $-\log(2\cosh J_{2})-\frac{1}{M}\log(1+(\tanh J_{2})^{M})$. We can substitute $a=\coth(2J_{2})$ in the left side of Eq. (4.12), and $\coth J_{2}=a+\sqrt{a^{2}-1}$ in the right side.
@@
We finally discuss the “quantum phase transition” of the quantum Ising model. The free energy $f^{\rm qu}(\beta,h)$ of the one-dimensional model is clearly analytic for all $\beta>0,h\in{\mathbb{R}}$ (and in a complex neighbourhood), but interesting behaviour can happen in the zero temperature limit. Namely, we consider the ground state energy
@@
around $h=-\frac{1}{2}$ and $h=\frac{1}{2}$. As is well-known, there are multiple ground states when $|h|<\frac{1}{2}$, that display long-range order; there is a single disordered ground state when $|h|>\frac{1}{2}$. More information about the quantum Ising model can be found in the recent works [8, 4, 3, 17, 25].

### v2

@@ Abstract.
We give a rigorous derivation of the free energy of (i) the classical Ising model on the triangular lattice with translation-invariant coupling constants, and (ii) the one-dimensional quantum Ising model. We use the method of Kac and Ward. The novel aspect is that the coupling constants may have negative signs. We describe the logarithmic singularity of the specific heat of the classical model and the validity of the Cimasoni–Duminil-Copin–Li formula for the critical temperature. We also discuss the quantum phase transition of the quantum model.
@@ 1. Introduction
After the work of Onsager and Kaufman, people found two alternate approaches: combinatorial and fermionic. The former was proposed in 1952 by Kac and Ward [14]; it was later extended by Kasteleyn who noted the connection with dimer systems [16] (see also Temperley and Fisher [28]). Potts [24] and Stephenson [26] used the Kac-Ward method on the triangular lattice, for the free energy and for correlation functions. The fermionic method was proposed in 1964 by Schultz, Mattis, and Lieb [25].
@@
The main goal of this article is to extend the Kac-Ward method to the case of (translation-invariant) coupling constants of arbitrary signs. We work on the triangular lattice, which is the simplest case of frustrated systems with translation-invariant coupling constants. We start with the Cimasoni extension of the Kac-Ward formula to “faithful projections” of non-planar graphs [5] (see also Aizenman and Warzel [2] for a clear exposition). We use it for the torus $\{1,\dots,L\}_{\rm per}\times\{1,\dots,M\}_{\rm per}$ with periodic boundary conditions. The main difficulties involve the non-planarity of the graph. We prove that these difficulties vanish in the limit $L\to\infty$ for fixed $M$. Then we can use the Fourier transform and we obtain the free energy formula for the infinite cylinder ×${\mathbb{Z}}\times\{1,\dots,M\}_{\rm per}$. The Onsager-Houtappel formula immediately follows by taking the limit $M\to\infty$.
@@
Our result for cylinders allows us to consider the one-dimensional quantum Ising model, whose free energy was first calculated in 1970 by Pfeuty [23]. We refer to [9, 4, 13, 7, 3, 19, 27] for recent studies. The quantum Ising model can be mapped to a 2D classical Ising model in the limit where the extra dimension becomes continuous. We also discuss the occurrence of a “quantum phase transition”.
@@ 2.1. The free energy
As $M\to\infty$ the cylinder free energy density converges to the two-dimensional free energy density:
@@
Figure 2. Plots of the free energy $\mathsf{f}(\beta)$ and its first and second derivatives for the translation-invariant triangular lattice ($J_{1}=J_{2}=J_{3}=1$). The second derivative has a logarithmic singularity at $\beta_{\rm c}=\frac{1}{4}\log 3=0.274...$.
@@ 2.2. Logarithmic singularity of the specific heat
We are looking for values of $\beta$ where ${\mathsf{f}}$ is not analytic. We show the well-known fact that the second derivative of ${\mathsf{f}}$ (which is related to the physical quantity called the specific heat) has a logarithmic singularity at a special value $\beta_{\rm c}$, called the critical point. This is illustrated in Fig. 2, which displays the free energy ${\mathsf{f}}(\beta)$ and its first and second derivatives in the case of the homogenous triangular lattice ($J_{1}=J_{2}=J_{3}=1$). By Theorem 2.1 (b) we have
@@
It is not hard to verify that the second condition holds true when $\sinh(2\beta_{\rm c}J_{i})+2\sinh(2\beta_{\rm c}J_{3})>0$ for $i=1,2$.
@@
We already know that ${\mathsf{f}}(\beta)$ is concave and therefore continuous. For $\beta\neq\beta_{\rm c}$, we have
@@
We have now verified that the only divergent term in (2.18) is the first one, and the divergence is logarithmic indeed. ∎
@@ 2.3. Case $J_{1}=J_{2}$
    If $J_{3}>-1$, there is a unique $\beta_{\rm c}$ such that ${\mathsf{f}}(\beta)$ is analytic in +${\mathbb{R}}_{+}\setminus\{\beta_{\rm c}\}$ and ${\mathsf{f}}^{\prime\prime}(\beta)$ has a logarithmic divergence at $\beta_{\rm c}$.
@@
The theorem is illustrated with the phase diagram of Fig. 3.
Figure 3. Phase diagram with $J_{1}=J_{2}=1$. The free energy is proved to lack analyticity at the line that separates the “ordered” and “uniqueness” phases. The separation line is the inverse critical temperature $\beta_{\rm c}=\beta_{\rm c}(J_{3})$; it is solution of the equation $\tanh\beta_{\rm c}=j^{-1}(J_{3})$ with $j$ defined in Eq. (2.26). For $J_{3}\geq 0$, the article [6] proves the existence of a unique infinite-volume Gibbs state for $\beta<\beta_{\rm c}$, and of several distinct Gibbs states for $\beta>\beta_{\rm c}$. For $J_{3}<0$, uniqueness is only proved for small $\beta$, and the existence of multiple Gibbs states is only proved for large $\beta$ (using the Pirogov-Sinai theory, see e.g. [8]).
@@
In order to make the connection to the free energy (2.12), we remark that
@@
Then $g(\beta)$ vanishes precisely when $a(\beta)$ does; indeed, $h(\beta;k_{1},k_{2})$ is nonnegative when the coupling constants are nonnegative, and its minimum is 0. Proposition 2.5 applies, establishing the singularity of the second derivative of the free energy.
We check in the proof below that the Cimasoni–Duminil-Copin–Li formula (2.20) holds whenever $J_{1}=J_{2}\geq 0$, and $J_{3}\in{\mathbb{R}}$ is allowed to be negative. One can also check that it does not hold if $J_{1}=J_{2}$ change signs; indeed, the free energy is the same due to symmetries (spin flips on a sublattice), but Eq. (2.20) is different and has different solutions.
In addition to the non-analyticity of the free energy, Cimasoni and Duminil-Copin prove that the phase transition involves a change of the number of infinite-volume Gibbs states: There is just one for $\beta\leq\beta_{\rm c}$ and more than one for $\beta>\beta_{\rm c}$. The proof relies on the GKS and FKG correlation inequalities, which hold for nonnegative coupling constants only. It would be interesting to extend this to the case of coupling constants with arbitrary signs.
@@ 3. The Kac-Ward identity
We rely on the extension of the Kac-Ward identity to “faithful projections” of non-planar graphs. It was proposed by Cimasoni [5] and used in [15, 2]. In order to accommodate negative weights we need two faithful projections for ${\mathbb{T}}_{L,M}$ with edges between nearest-neighbours. The graphs are $G_{1}$ and $G_{2}$ and they are illustrated in Fig. 4. Here is a full description of the left graph:
@@
  • There are edges represented by “handles” (continuous curves with winding number $-1$) between $(L,j)$ and $(1,j)$ for $1\leq j\leq M$; between $(L,j)$ and $(1,j+1)$ for $11\leq j\leq M-$ 1; between $(i,M)$ and $(i,1)$ for $1\leq i\leq L$; between $(i,M)$ and $(i+1,1)$ for $11\leq i\leq L-$ 1.
@@
It is worth noting that the right side of (3.3) is a multinomial in $(W_{e,e})$, something that is not apparent in the left side — there are remarkable cancellations indeed. This allows [2] to prove the identity for small $(W_{e,e})$; the extension to larger values is automatic. The determinant cannot be negative and the sign of the square root cannot change.
We define the matrix $K^{\scriptscriptstyle(2)}$ as in (3.1) but $\measuredangle_{2}(e,e^{\prime})$ and $\measuredangle_{2}(e)$ are the corresponding angles on the the faithfull projection $G_{2}$. Analogously, we define $n^{\scriptscriptstyle(2)}_{0}(\Gamma)$ for this projection.
The connection with the Ising model is through the high-temperature expansion, see e.g. [8, Section 3.7.3]. The partition function (2.4) is equal to
@@
$\sqrt{\det(1-K^{\scriptscriptstyle(1)}W)}+\sqrt{\det(1-K^{\scriptscriptstyle(2)}W)}=2\widetilde{Z}_{L,M}\Bigl{(}1-\big{\langle}1_{n_{\rm h}(\Gamma)\,\rm{odd}}\big{\rangle}_{L,M}\Bigr{)}.$
@@
We condition on the horizontal handles (including possibly the self-crossing ones). We denote by ${\mathfrak{h}}$ the set of handles that connect sites in the leftmost and rightmost columns:
@@
Then we define the support ${\rm{supp\,}}_{1}{\mathfrak{h}}$, resp.  ${\rm{supp\,}}_{L}{\mathfrak{h}}$, to be the set of vertices of the form $(1,j_{i})$, resp.  $(L,j_{i}^{\prime})$, that appear an odd number of times in ${\mathfrak{h}}$. Let ${\rm{supp\,}}{\mathfrak{h}}={\rm{supp\,}}_{1}{\mathfrak{h}}\cup{\rm{supp\,}}_{L}{\mathfrak{h}}$. We let ~$\tilde{\mathcal{E}}_{L,M}$ be the set of edges of the cylinder (not the torus) $\{1,\dots,L\}\times{\mathbb{T}}_{M}$. With $1_{\mathfrak{h}}=1_{\mathfrak{h}}(\Gamma)$ the indicator function that the random graph $\Gamma$ has set of handles ${\mathfrak{h}}$, we have
$\begin{split}\big{\langle}1_{n_{\rm h}(\Gamma)\,\rm{odd}}\big{\rangle}_{L,M}&=\sum_{|{\mathfrak{h}}|\;{\rm odd}}\langle 1_{\mathfrak{h}}\rangle_{L,M}\\ &=\sum_{|{\mathfrak{h}}|\;{\rm odd}}\biggl{(}\prod_{i=1}^{k}\tanh J_{(1,j_{i}),(L,j_{i}^{\prime})}\biggr{)}\frac{1}{\widetilde{Z}_{L,M}}\sum_{\Gamma\subset\tilde{\mathcal{E}}_{L,M}:\partial\Gamma={\rm{supp\,}}{\mathfrak{h}}}w(\Gamma).\end{split}$ (3.9)
@@
One can expand the determinants as products of directed loops as in [2, Theorem 3.2]. Let $\gamma=(e_{1},\dots,e_{k})$ be a directed loop with $\ell$ handles (the self-crossing handle between $(L.M)$ and $(1,1)$ is counted twice). We have
$\prod_{i=1}^{k}K^{\scriptscriptstyle(1)}_{e_{i},e_{i+1}}=(-1)^{\ell}\,\prod_{i=1}^{k}\widetilde{K}_{e_{i},e_{i+1}},\qquad\prod_{i=1}^{k}\widetilde{W}^{\scriptscriptstyle(1)}_{e_{i},e_{i}}=(-1)^{\ell}\,\prod_{i=1}^{k}W_{e_{i},e_{i}}.$ (3.17)
Then each loop gives the same contribution in $\det(1-K^{\scriptscriptstyle(1)}W)$ and in $\det(1-\widetilde{K}\widetilde{W}^{\scriptscriptstyle(1)})$. The argument for $\det(1-K^{\scriptscriptstyle(2)}W)$ is the same, counting only vertical and oblique handles between sites $(i,M)$ and $(j,1)$, $1\leq i,j,\leq L$. ∎
@@
where $t_{i}=\tanh{J_{i}}$, $i=1,2,3$. Then
@@
(a) By Eq. (3.3) and Lemma 3.3, we obtain that both square roots of the above determinants are real. (b) This is a consequence of Lemma 3.4; taking the logarithm we obtain Riemann sums. ∎
@@ 4. The 1D quantum Ising model
One application of the cylinder formula of Theorem 2.1 (a) deals with the one-dimensional quantum Ising model. It is well-known that it can be mapped to a classical model in $1+1$ dimensions, the extra dimension being the continuous interval $[0,\beta]$ with periodic boundary conditions. A phase transition is only possible when both dimensions are infinite, which necessitates taking the limit of zero-temperature $\beta\to\infty$. The free energy of the quantum Ising model was first computed by Pfeuty [23] using the fermionic method of [25]. The results of this section are not new, but the Kac-Ward approach may have more appeal to some readers.
@@
Notice the division by $\beta$, which allows to get the ground state energy by taking the limit $\beta\to\infty$.
@@
$\begin{split}{\operatorname{Tr\,}}\,{\rm e}^{-\beta H_{L}}\,&=\lim_{n\to\infty}{\operatorname{Tr\,}}\biggl{(}\,{\rm e}^{\frac{\beta}{n}\sum_{i=1}^{L}S_{i}^{\scriptscriptstyle(3)}S_{i+1}^{\scriptscriptstyle(3)}}\,\prod_{i=1}^{L}\bigl{(}1+\tfrac{\beta h}{n}S_{i}^{\scriptscriptstyle(1)}\bigr{)}\biggr{)}^{n}\\ &=\lim_{n\to\infty}\sum_{\sigma^{\scriptscriptstyle(1)},\dots,\sigma^{\scriptscriptstyle(n)}}\exp\biggl{\{}\frac{\beta}{4n}\sum_{i=1}^{L}\sum_{k=1}^{n}\sigma_{i}^{\scriptscriptstyle(k)}\sigma_{i+1}^{\scriptscriptstyle(k)}\biggr{\}}\prod_{i=1}^{L}\prod_{k=1}^{n}\langle\sigma_{i}^{\scriptscriptstyle(k)}|\bigl{(}1+\tfrac{\beta h}{n}S^{\scriptscriptstyle(1)}\bigr{)}|\sigma_{i}^{\scriptscriptstyle(k+1)}\rangle.\end{split}$ (4.5)
@@
    For fixed $n$ the limit $L\to\infty$ of $f_{L,n}^{\rm qu}(\beta,h)$ exists (and is denoted $f_{\infty,n}^{\rm qu}(\beta,h)$).
@@
The first equality in (b) is clear. For the second equality we use the following estimates, which again follow from estimates on the classical partition function:
@@
$f_{L,n}^{\rm qu}(\beta,h)+\tfrac{1}{2L}\geq f_{\infty,n}^{\rm qu}(\beta,h)\geq f_{L,n}^{\rm qu}(\beta,h)-\tfrac{1}{2L}.$ (4.9)
@@
This holds for any ${\varepsilon}>0$ provided $n$ is large enough. This proves the second identity in (b). ∎
@@
It can be obtained by taking the limit $J_{1}\to 0$ in Theorem 2.1 (a), as the expression converges to the free energy of the 1D Ising model in ${\mathbb{T}}_{M}$. The latter is easily calculated with the 1D transfer matrices, yielding $-\log(2\cosh J_{2})-\frac{1}{M}\log(1+\tanh^{M}J_{2})$. We can substitute $a=\coth(2J_{2})$ in the left side of Eq. (4.12), and $\coth J_{2}=a+\sqrt{a^{2}-1}$ in the right side.
@@
We finally discuss the “quantum phase transition” of the quantum Ising model. The free energy $f^{\rm qu}(\beta,h)$ of the one-dimensional model is clearly analytic for all $\beta>0,h\in{\mathbb{R}}$ (and in a complex neighbourhood), but interesting behaviour can happen in the zero-temperature limit. Namely, we consider the ground state energy
@@
around $h=-\frac{1}{2}$ and $h=\frac{1}{2}$. As is well-known, there are multiple ground states when $|h|<\frac{1}{2}$, that display long-range order; there is a single disordered ground state when $|h|>\frac{1}{2}$. More information about the quantum Ising model can be found in the recent works [9, 4, 13, 7, 3, 19, 27].